OpenStudy (anonymous):

im working with the unit circle to find values of tan between (0,360 dgrees) I factored this problem and have tan(theta)=3,-3 How do I convert to cos and sin so I can locate the correct angles on the unit circle?

OpenStudy (anonymous):

i use identity tan = sin/cos but I only have my tan value

imqwerty (imqwerty):

so u want to find sin and cos and u have tan(theta)=3, -3 given right?

OpenStudy (anonymous):

ya i factored and got those 2 results i need angles between 0,360 but i need to convert so i cn use the unit circle

OpenStudy (anonymous):

but I forgot how i think

imqwerty (imqwerty):

u shuld try this identity-\[\sec^2 \theta=1+\tan^2 \theta\] from here u will get 2 values of sec sec =1/cos so u get 2 values of cos one will be positive the other will be negative ok till here?

OpenStudy (anonymous):

oooo that was it ok! wow

OpenStudy (anonymous):

Thanks a lot!

OpenStudy (anonymous):

i should probably memorize these identities somehow so i dont have to keep eyeballing it during the test

OpenStudy (anonymous):

im just really bad at trig, its all so confusing

OpenStudy (anonymous):

im good from there though thanks much

imqwerty (imqwerty):

yes u will get the value of sec^2 (theta) and we know that sec^2(theta)=1/cos^2(theta) so u have 2 values of cos then u use this identity-\[\sin^2 \theta+\cos^2\theta=1\] u have the value of cos^2(theta) u can get sin^2(theta) again there will be 2 sin(theta) one positive one negative)